# Eaman Jahani

### v1

@@ Abstract
In an online experiment with N = 1,891 participants, we collected and analyzed over 18,000 prompts to explore how the importance of prompting will change as the capabilities of generative AI models continue to improve. Each participant in our experiment was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. Participants were then asked to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. More specifically, despite being blind to the model they were assigned, participants assigned to DALL-E 3 wrote longer prompts that were more semantically similar to each other and contained a greater number of descriptive words. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Taken together, our results suggest that as models continue to progress, people will continue to adapt their prompts to take advantage of new models’ capabilities.
@@
In this paper, we bring experimental evidence to bear on this debate by presenting results from a pre-registered online lab experiment with $N=1,891{}$ participants recruited on Prolific. [^1] This paper presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI. In the experiment, participants were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) DALL-E 3, and (3) a version of DALL-E 3 with automatic LLM-based prompt revisions (henceforth referred to as “DALL-E 3 with revision”). These prompt revisions were not made visible to participants during the experiment. [^2] Automatic prompt revision is the default behavior of the DALL-E 3 API endpoint. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions. For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
@@ Results
The majority of our analyses focus on the DALL-E 2 and DALL-E 3 treatment arms, as they are more directly comparable. We first estimate the effect of access to a more advanced generative model on task performance by evaluating how similar each generated image is to the relevant target image. [^3] All main text analyses use ”replay” data from Figure 2 for consistency and to avoid model drift issues. All results hold with the “original” data. See the SI for details. To construct a quantitative measure of image similarity, we generated CLIP embedding vectors (Radford et al., 2021) for each participant-generated image and the relevant target image and then calculated the cosine similarity ($CoSim$) between those vectors (see SI for further details on how we quantify image similarity and account for the stochasticity of model output). Unsurprisingly, we find that providing access to a more advanced generative model improved task performance. Participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.20$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model). Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in similarity to the target image. We also find that the treatment effect increased slightly as participants made successive attempts to replicate the target image (See Figure 1B; $\beta=0.0008$, $p=0.042$). In other words, the performance gap widened between DALL-E 2 and DALL-E 3 as participants gained more experience with the respective models, potentially due to changes in prompting.
@@
To investigate how participants’ prompts changed in response to the superior model’s capabilities, we conducted an exploratory analysis of the contents of participant prompts and how those prompts evolved as participants iterated. We find that participants assigned the superior model wrote prompts that were 24% longer ($\Delta Words=6.6$, $p<10^{-4}$), and that the difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-7}$). Despite differing in length, prompts written for both models contained similar proportions of nouns and adjectives (48% superior vs. 49% inferior; $p$ = $0.148$). This suggests that prompts submitted to DALL-E 3 were longer because they conveyed additional descriptive information, not because they were padded with superfluous words. We also find that participants blindly assigned to DALL-E 3 wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0166$, $p=0.02$) and in the aggregate ($\beta=0.0169$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses). These findings indicate that exposure to different models did in fact induce differences in prompting, and that these differences grew over successive replication attempts.
Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects. The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line). Effects are shown relative to a baseline of prompts written for and replayed on the inferior model. Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
@@
To understand the extent to which the ATE we observe is attributable to these behavioral differences, as opposed to DALL-E 3’s more advanced capabilities, we conducted another exploratory analysis in which we regenerated images with (or “replayed”) all participant prompts on both DALL-E 2 and DALL-E 3, regardless of the model the participant used during the experiment. This allows us to decompose the ATE into two components: the “model effect,” i.e., the average increase in similarity to the target when replaying the prompts written by inferior model participants on the superior model, and the “prompting effect,” i.e., the difference between the average similarity to the target of all prompts generated by superior model participants and the average similarity to the target of all prompts generated by inferior model participants, both replayed on the superior model. Figure 2 shows that the model effect (red) accounts for 53% of the ATE ($\Delta CoSim=0.00864$, $p<10^{-8}$), whereas the prompting effect (blue) accounts for 47% of the ATE ($\Delta CoSim=0.00769$, $p=0.0066$). Thus, nearly half of the ATE is driven by endogenous changes in prompting strategy that emerged after exposure to DALL-E 3. The rightmost point in Figure 2 shows that these endogenous changes in prompting took advantage of the unique capabilities of the superior model; when replaying both sets of prompts on the inferior model, the difference in performance is not statistically significant and is close to zero ($\Delta CoSim=0.0018$, $p=0.52$). In other words, DALL-E 3 was more capable than DALL-E 2 at effectively rendering the additional information contained in the longer and more descriptive prompts (Betker et al., 2023), and participants adapted their prompting behavior accordingly.
@@
One might think that the need to adapt one’s prompts to the capabilities of different generative models is obviated by LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023). We use data from our third treatment arm—DALL-E 3 with revision—to test this hypothesis. We find that although participants assigned to DALL-E 3 with revisions performed better than those assigned to DALL-E 2 ($\Delta CoSim=0.0068$; $p=0.013$), prompt revision reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [22%, 94%]). Importantly, the positive impact of access to DALL-E 3 with revision was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). In other words, prompt revision actually caused participants to perform worse in our context. This finding suggests that as currently implemented, AI-assisted prompt revisions are not a panacea, and can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
@@ Discussion
In summary, we find that the progression of generative AI models from DALL-E 2 to DALL-E 3 caused performance to increase, not only because of DALL-E 3’s superior capabilities (the model effect), but also because of changes in prompting behavior that took advantage of those capabilities (the prompting effect). More specifically, participants using DALL-E 3 wrote longer prompts that contained more descriptive information, and DALL-E 3 was able to render this additional information more effectively than DALL-E 2. These changes in prompting emerged despite the fact that participants were unaware of which model they were assigned. Although automated LLM-revision features are designed to and have been shown to help people generate higher-quality content (Betker et al., 2023; Li et al., 2024), we also find that LLM-based prompt revision failed to further improve performance; instead, automatic prompt revision reduced the benefits of using DALL-E 3. Overall, we find that shifting to a superior model, whether or not accompanied by LLM-based revisions, did not eliminate the need for prompting.
@@ A.2 Randomization
We randomized participants across two dimensions: the target image and the text-to-image generative AI model that participants had access to. We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target image-model cells. We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test for homogeneity of proportions across all cells. With $\chi^{2}=0.374,\;df=28$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
@@ A.2.1 Generative Models
In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process. In the DALL-E 3 (Verbatim) treatment arm, we prepend a string instructing the GPT-4 model to not modify the participant’s prompt before passing it forward to DALL-E 3. This string is never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint. [^4] See here and here for online documentation. We modified the recommended preflix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
@@ A.2.2 Target Images
We randomly assigned participants to 1 of 15 target images. The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and photography. We chose these to represent the use cases suggested by the prompt categories on https://promptbase.com/, a leading marketplace for image generation prompts. The images vary in color, style, content, and complexity within and across categories. These images can be found online linked to the pre-registration document: https://osf.io/ejbtp. As we discuss in Section F, performance, and variability of performance varied substantially across images. In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
@@ B.4 Sample Construction
  • We next removed participants from our sample if they failed to submit at least 10 prompts or if they did not make a “good-faith” effort to complete the task as requested. Both of these exclusion criteria were pre-registered. A good-faith effort was not made if a participant submitted the same prompt at least five times in a row at any point during the task. These exclusion criteria were also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” To avoid reward hacking, we did not specify the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,899 participants.
@@
    • While analyzing our data, we found that our sample contained a number of prompts that were not “good-faith efforts” to complete our task, despite the use of the exclusion criteria described above. Thus, we used the following process to identify and remove additional “low-effort” prompts. First, we generated embeddings for each prompt using OpenAI’s text-embedding-3-small model. We then calculated the mean embedding for each target image. Next, we calculated the Euclidean distance between each prompt’s embedding vector and the mean embedding vector for prompts corresponding to the focal prompt’s assigned target image. Finally, we removed the 2.5% of prompts that were most distant from the mean image-level prompt embedding vector. We observed that the prompts that were dropped were often off-topic and likely were not good-faith efforts to complete our task. This led to the removal of 481 prompts across all three treatment arms, and 338 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms. Our results are robust to the inclusion of these filtered prompts.
  • Our final sample included 1,891 participants and 18,152 prompts.
@@ B.5.1 Image Similarity
  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al., 2023). The authors claim that relative to something like CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.

### v2

@@ Abstract
In an online experiment with N = 1,893 participants, we collected and analyzed over 18,000 prompts and over 300,000 images to explore how the importance of prompting will change as the capabilities of generative AI models continue to improve. Each participant in our experiment was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. Participants were then asked to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. More specifically, despite being blind to the model they were assigned, participants assigned to DALL-E 3 wrote longer prompts that were more semantically similar to each other and contained a greater number of descriptive words. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Taken together, our results suggest that as models continue to progress, people will continue to adapt their prompts to take advantage of new models’ capabilities.
@@
In this paper, we bring experimental evidence to bear on this debate by presenting results from a pre-registered online lab experiment with $N=1,893{}$ participants recruited on Prolific. [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI. In the experiment, participants were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”). The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment. [^2] Automatic prompt revision is the default behavior of the DALL-E 3 API endpoint. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions. For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
@@ Results
The majority of our analyses focus on the DALL-E 2 and DALL-E 3 treatment arms, as they are more directly comparable. We first estimate the effect of access to a more advanced generative model on task performance. [^3] All main text analyses use ”replay” data from Figure 2 for consistency and to avoid model drift issues. All results hold with the “original” data. See the SI for details. Given the stochastic nature of generative AI models, our measure of task performance is the expected similarity of images generated from each prompt to the relevant target image. To construct a quantitative measure of this expected image similarity for each prompt, we first used each participant-generated prompt to produce 10 images. We then computed CLIP embedding vectors (Radford et al., 2021) for both target images and prompt-generated images. For each participant prompt and its corresponding target image, we then calculated the average cosine similarity ($CoSim$) between the embedding vectors of the randomly sampled generated images and the target image vector (see SI for further details on how we quantify expected image similarity). Unsurprisingly, we find that providing access to a more advanced generative model improved task performance. Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2. Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in similarity to the target image. We also find that the treatment effect increased slightly as participants made successive attempts to replicate the target image (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression). In other words, the performance gap widened between DALL-E 2 and DALL-E 3 as participants gained more experience with the respective models, potentially due to changes in prompting.
@@
To investigate how participants’ prompts changed in response to DALL-E 3’s more advanced capabilities, we conducted an exploratory analysis of the contents of participant prompts and how those prompts evolved as participants iterated. We find that participants assigned DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and that the difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$). Despite differing in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs. 49% DALL-E 2; $p$ = $0.215$). This suggests that prompts submitted to DALL-E 3 were longer because they conveyed additional descriptive information, not because they were padded with superfluous words. We also find that participants blindly assigned to DALL-E 3 wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses). These findings indicate that exposure to different models did in fact induce differences in prompting, and that these differences grew over successive replication attempts.
@@
To understand the extent to which the ATE we observe is attributable to these behavioral differences, as opposed to DALL-E 3’s more advanced capabilities, we conducted another exploratory analysis in which we regenerated images with (or “replayed”) all participant prompts on both DALL-E 2 and DALL-E 3, regardless of the model the participant used during the experiment. This allows us to decompose the ATE into two components: the “model effect,” i.e., the average increase in similarity to the target when replaying the prompts written by DALL-E 2-assigned participants on DALL-E 3, and the “prompting effect,” i.e., the difference between the average similarity to the target of all prompts generated by DALL-E 3-assigned participants and the average similarity to the target of all prompts generated by DALL-E 2-assigned participants, both replayed on DALL-E 3. Figure 2 shows that the model effect (red) accounts for 51% of the ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$), whereas the prompting effect (blue) accounts for 48% of the ATE ($\Delta CoSim=0.00788$, $p=0.024$). Thus, nearly half of the ATE is driven by endogenous changes in prompting strategy that emerged after exposure to DALL-E 3. The rightmost point in Figure 2 shows that these endogenous changes in prompting took advantage of the unique capabilities of DALL-E 3; when replaying both sets of prompts on DALL-E 2, the difference in performance is not statistically significant and is close to zero ($\Delta CoSim=0.0020$, $p=0.56$). In other words, DALL-E 3 was more capable than DALL-E 2 at effectively rendering the additional information contained in the longer and more descriptive prompts (Betker et al., 2023), and participants adapted their prompting behavior accordingly.
@@
One might think that the need to adapt one’s prompts to the capabilities of different generative models can be obviated by LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023). We use data from our third treatment arm—DALL-E 3 with revision—to test this hypothesis. We find that although participants assigned to DALL-E 3 with revisions performed better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$), prompt revision reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]). Importantly, the positive impact of access to DALL-E 3 with revision was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). In other words, prompt revision actually caused participants to perform worse in our context. This finding suggests that as currently implemented, AI-assisted prompt revisions are not a panacea, and can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
@@ Discussion
Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects. The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line). Effects are shown relative to a baseline of prompts written for and replayed on DALL-E 2. Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
In summary, we find that the progression of generative AI models from DALL-E 2 to DALL-E 3 caused performance on our image replication task to increase. This was not only because of DALL-E 3’s superior capabilities (the model effect), but also because of changes in prompting behavior that took advantage of those capabilities (the prompting effect). More specifically, participants using DALL-E 3 wrote longer prompts that contained more descriptive information, and DALL-E 3 was able to render that additional information more effectively than DALL-E 2. These changes in prompting emerged despite the fact that participants were unaware of which model they were assigned. Although automated LLM-revision features are designed to and have been shown to help people generate higher-quality content (Betker et al., 2023; Li et al., 2024), we also find that LLM-based prompt revision failed to further improve performance; instead, automatic prompt revision reduced the benefits of using DALL-E 3. Overall, we find that shifting to a more capable model, whether or not accompanied by automatic LLM revisions, did not eliminate the need for prompting.
@@ A.2 Randomization
We randomized participants across two dimensions: the target image and the text-to-image generative AI model that participants had access to. We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target image-model cells. We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test across all cells. With $\chi^{2}=7.056,\;df=44$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
@@ A.2.1 Generative Models
In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process. In the DALL-E 3 (Verbatim) treatment arm, we prepend a string instructing the GPT-4 model to not modify the participant’s prompt before passing it forward to DALL-E 3. This string is never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint. [^4] See here and here for online documentation. We modified the recommended prefix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
@@ A.2.2 Target Images
We randomly assigned participants to 1 of 15 target images. The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and architectural photography. We chose these to represent the use cases suggested by the prompt categories on https://promptbase.com/, a leading marketplace for image generation prompts. The images vary in color, style, content, and complexity within and across categories. These images can be found online linked to the pre-registration document: https://osf.io/ejbtp. As we discuss in Section F, performance, and variability of performance varied substantially across images. In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
@@ B.4 Sample Construction
  • We next removed participants from our sample if they failed to submit at least 10 prompts or if a participant submitted the same prompt at least five times in a row at any point during the task. Both of these exclusion criteria were pre-registered. These exclusion criteria were also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” To avoid reward hacking, we did not specify the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,899 participants.
@@
  • Our final sample included 1,893 participants and 18,560 prompts.
    B.4.1 “Off-Topic” Robustness Check
    While analyzing our data, we found that our sample contained a number of “off-topic” prompts that did not seem related to the task. As a robustness check on our main results, we used the following process to systematically identify and remove “off-topic” prompts. First, we generated embeddings for each prompt using OpenAI’s text-embedding-3-small model. We then calculated the mean embedding for each target image. Next, we calculated the Euclidean distance between each prompt’s embedding vector and the mean embedding vector for prompts corresponding to the focal prompt’s assigned target image. Finally, we removed the 2.5% of prompts that were most distant from the mean image-level prompt embedding vector. This led to the removal of 481 prompts across all three treatment arms, and 338 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms. All of our main text results are robust to the exclusion of these “off-topic” prompts.
@@ B.5.1 Image Similarity
  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al., 2023). The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.